A Tanzanian truck driver traveling at 22.5m/s when approaching a primary school decelerates at 3.0m/s2 before he stopped.

(a) How much time does it for the truck driver to stop?

(b) How far does he travel while stopping?


1
Expert's answer
2021-02-25T18:32:48-0500

(a)

"v=v_0+at,""t=\dfrac{v-v_0}{a}=\dfrac{0-22.5\ \dfrac{m}{s}}{-3.0\ \dfrac{m}{s^2}}=7.5\ s."

(b)

"d=\dfrac{1}{2}(v_0+v)t,""d=\dfrac{1}{2}\cdot(22.5\ \dfrac{m}{a}+0)\cdot7.5\ s=84.4\ m."